OpenStudy (ajoy31):

How do I simplify (-3-i)/(4+8i)? i being the square root of -1.

OpenStudy (anonymous):

Distribute, you get -12 - 24i - 4i -8i^2 = -12 -28i +8 = 20-28i NOTE: i^2 = -1, so -8i^2 = +8

OpenStudy (ajoy31):

So I would change the division to multiplication...?

myininaya (myininaya):

I think @easyaspi314 since the operation as multiplication while it is division. Multiply top and bottom by bottom's conjugate.

myininaya (myininaya):

seen*

OpenStudy (anonymous):

presumably the question means "divide, write the answer in standard form" correct?

OpenStudy (anonymous):

\[\frac{-3-i}{4+8i}\times \frac{4-8i}{4-8i}=\frac{(-3-i)(4-8i)}{4^2+8^2}\]

OpenStudy (anonymous):

exactly